OpenStudy (anonymous):

For which equation is x=-1/2 a solution? A. 7 – 4x = –9 B. 9x + 2 = 7 C. 6 + 8x = –10 D. 10x – 3 = –8

OpenStudy (anonymous):

7-4x=-9 -4x=-16 x=4 A aint it

OpenStudy (anonymous):

Plug -1/2 into each equation until you can find the right answer

OpenStudy (anonymous):

9x+2=7 9x=5 x=5/9 B aint it

OpenStudy (anonymous):

6+8x=-10 8x=-16 x=-2 C aint it

OpenStudy (anonymous):

10x-3=-8 10x=-5 x=-1/2 D is the answer